# Can LLM “Self-report”?: Evaluating the Validity of Self-report Scales in Measuring Personality Design in LLM-based Chatbots

### v1

@@ 6 Conclusion
This paper demonstrates the limitations of self-reported personality assessments in evaluating the personality design of LLM-based chatbots. Our findings reveal the discrepancy between chatbots’ self-reported personality scores and human task-based perceptions, suggesting that self-assessments may not accurately capture how chatbots are perceived in real-world interactions. Additionally, our analysis of predictive validity indicates that self-reported personality scales do not align with interaction quality. These results highlight the need for evaluation methods that capture chatbot personality in task-driven, interactive scenarios.
Limitations
Several limitations exist in the current study. First, there may be bias in the choice of psychometric tests. Although we sought to minimize this by using personality assessments of different lengths for self-reported data, the human-perceived personality was measured with only one questionnaire due to time constraints. Future research could address this by incorporating a broader range of personality tests and developing assessments specifically tailored for LLMs to achieve more accurate measurements.
Second, our findings are based on a single chatbot personality setting method and may not generalize well to others, highlighting the need for further investigation with different approaches. Additionally, the chatbot was designed with a focus on task completion, mirroring real-world applications to ensure ecological validity. As a result, the task settings may have constrained the variability in perceived personality. Moreover, the evaluation was conducted on five common chatbot tasks, which may not capture the full spectrum of user interactions. Expanding the task set in future studies could provide a more comprehensive view of task-based chatbot personality assessment.
Third, while our study only focuses on GPT-4o OpenAI (2024), we expect the same experimental setup could be extended to other LLMs. The extent to which our findings generalize to a broader range of models remains an open question for future research. Furthermore, while LLMs perform well on benchmarks across multiple languages, this study evaluated the model using English psychometric tests and English-speaking participants. Future research could explore personality dimensions within culture-specific contexts to provide more diverse insights on chatbot design.
@@
Our human study was reviewed and granted Exempt status by the Institutional Review Board (IRB). Participants are paid at the rate of \$12 per hour through the Prolific platform. We ensure transparency by explicitly stating the study’s purpose, requirements, and payments in the task instructions and obtaining informed consent from participants before the study begins. Additionally, we manually verify that all collected data are free of sensitive or personally identifiable information. All prompts used in this study are included in the Appendix. The codes for collecting self-report personality scores and the transcripts of human-chatbot interaction are publically accessible  [^2] https://github.com/isle-dev/self-report. The cost of running the self-report experiments with the batch API of GPT-4o is \$37.41.
@@ Appendix A Chatbot Prompt Format
The role defines the specific role the chatbot simulates across five tasks, such as a supportive companion or an educational guide. The level describes the extent to which the chosen domain is represented. The domain specifies the personality dimension selected from the Big Five model John et al. (1999). The profile describes the chatbot’s personality using five key adjectives, each paired with a qualifier, as outlined in section 2.1.1. The objectives summarize the chatbot’s goals within the given task, such as offering personalized social support or explaining complex concepts. Finally, the additional_info section includes any extra specific focus areas for the task.
@@
Appendix B Self-reported Personality Prompt Format
A complete prompt format for self-reported personality BFI-2-XS Soto and John (2017b) and BFI-2 Soto and John (2017a) is as follows, where the personality_description is the same as the task-based personality description introduced in Section 2.1.1 and test_item represents each item in the psychometric tests.
@@ Appendix B Self-reported Personality Prompt Format
For the IPIP-NEO-120 Johnson (2014b) questionnaire, the response scale is modified as follows: 1 = very inaccurate, 2 = moderately inaccurate, 3 = neither accurate nor inaccurate, 4 = moderately accurate, and 5 = very accurate.
@@ Appendix E Dataset Statistics
In this study, we configured chatbot personalities across each Big Five domain John et al. (1999) by randomly sampling adjectives at high or low marker levels, generating ten distinct profiles per domain for each level. Each participant engaged with one chatbot in one of five predefined task settings and then completed the BFI-2-XS questionnaire Soto and John (2017b). We collected a total of 500 valid conversation transcripts and corresponding assessments, with each interaction averaging 8 turns. Table 9 summarizes the statistics of the human evaluation dataset.
Domain Personality Marker Level Task Total
5 10 High, 10 Low 5 500
@@ Appendix G Variance in Human-perceived Personality
Task Ext Agr Con Neu Ope
@@
Public Service 1.993 7.028 1.541 2.740 3.500
Social Support 1.189 2.724 1.009 4.287 3.517
@@
Table 10: F-value for human-perceived personality scores between high and low personality settings across tasks.
Note: n = 100 for each task. Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
@@
Table 10 presents the F-values for human-perceived personality scores under high and low personality settings across five tasks. These values indicate the ratio of variances in personality perceptions between the two settings for each domain. The table shows that variance between the high and low settings is more pronounced than within each group. Notably, this effect is most prominent in certain tasks and personality domains, such as agreeableness and conscientiousness.
Task Ext Agr Con Neu Ope
@@
Public Service 0.42 0.58 0.43 0.26 0.24
Social Support 0.08 0.63 -0.01 0.09 0.19
@@
Table 11: Correlation analysis between self-reported personality with BFI-2-XS and human-perceived personality scores across tasks.
Note: n = 500 in total. Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
Task Ext Agr Con Neu Ope
@@
Public Service 0.40 0.59 0.47 0.31 0.22
Social Support 0.02 0.64 -0.01 0.11 0.14
@@
Guided Learning 0.05 0.53 -0.03 0.30 0.08
Table 12: Correlation analysis between self-reported personality with BFI-2 and human-perceived personality scores across tasks.
Note: n = 500 in total. Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
Task Ext Agr Con Neu Ope
@@
Public Service 0.37 0.57 0.42 0.33 0.24
Social Support -0.02 0.48 -0.07 0.18 0.12
@@
Table 13: Correlation analysis between self-reported personality with IPIP-NEO-120 and human-perceived personality scores across tasks.
Note: n = 500 in total. Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
Appendix H Correlation Analysis
Table 11, 12 and 13 present the correlation analysis between self-reported and human-perceived personality scores using the BFI-2-XS, BFI-2, and IPIP-NEO-120 questionnaires, respectively. All three tables show a weak correlation between human-perceived personality and those evaluated via standard tests.

### v2

@@ 6 Conclusion
This paper highlights the limitations of chatbot’s self-report personality assessments in evaluating LLM-based chatbot personality design. Our findings reveal the discrepancy between chatbots’ self-repord personality scores and human task-based perceptions, suggesting that self-assessments may not accurately capture how chatbots are perceived in real-world interactions. Additionally, our analysis of predictive validity shows that self-report personality scales do not align with interaction quality. These results highlight the need for evaluation methods that capture chatbot personality in task-driven, interactive scenarios.
@@
Our human study was reviewed and granted Exempt status by the Institutional Review Board (IRB). Participants are paid at the rate of \$12 per hour through the Prolific platform. We ensure transparency by explicitly stating the study’s purpose, requirements, and payments in the task instructions and obtaining informed consent from participants before the study begins. Additionally, we manually verify that all collected data are free of sensitive or personally identifiable information.
Reproducibility Statement
All prompts used in this study are included in the Appendix. The code for collecting self-report personality scores and the human study data are publicly accessible [^2] https://github.com/isle-dev/self-report. The cost of running the self-report experiments using the GPT-4o batch API was \$37.41, while the cost of conducting the human study, including platform fees, was \$1091.57.
@@ Appendix A Chatbot Prompt Format
The role defines the specific role the chatbot simulates across five tasks, such as a supportive companion or an educational guide. The level describes the extent to which the chosen domain is represented. The domain specifies the personality dimension selected from the Big Five model (John et al., 1999). The profile describes the chatbot’s personality using five key adjectives, each paired with a qualifier, as outlined in section 3.1.1. The objectives summarize the chatbot’s goals within the given task, such as offering personalized social support or explaining complex concepts. Finally, the additional_info section includes any extra specific focus areas for the task.
@@
Appendix B Self-report Personality Prompt Format
A complete prompt format for self-report personality BFI-2-XS (Soto & John, 2017b) and BFI-2 (Soto & John, 2017a) is as follows, where the personality_description is the same as the task-based personality description introduced in Section 3.1.1 and test_item represents each item in the psychometric tests.
@@ Appendix B Self-report Personality Prompt Format
For the IPIP-NEO-120 (Johnson, 2014b) questionnaire, the response scale is modified as follows: 1 = “very inaccurate”, 2 = “moderately inaccurate”, 3 = “neither accurate nor inaccurate”, 4 = “moderately accurate”, and 5 = “very accurate”.
@@ Appendix E Dataset Statistics
In this study, we configured chatbot personalities across each Big Five domain (John et al., 1999) by randomly sampling adjectives at high or low marker levels, generating ten distinct profiles per domain for each level. Each participant engaged with one chatbot in one of five predefined task settings and then completed the BFI-2-XS questionnaire (Soto & John, 2017b), yielding 500 valid conversation transcripts (i.e., 10 chatbots × 2 levels × 5 domains × 5 tasks). We collected these transcripts and corresponding assessments, with each interaction averaging 8 to 9 turns. Table 8 summarizes the statistics of the human evaluation dataset.
Adjective Combination Marker Level Domain Task Total
10 2 5 5 500
@@ Appendix G Variance in Human-perceived Personality
Evaluation Task EXT AGR CON NEU OPE
Job Interview 1.089 2.759 2.949 4.413 3.389
Public Service 1.080 175.448 6.695 7.560 27.331
Self-report Social Support 1.076 126.509 6.869 6.172 1.463
Travel Planning 3.752 8.167 NA 7.490 1.256
Guided Learning 3.023 6.061 1.199 50.017 4.677
@@
Public Service 1.993 7.028 1.461 2.740 4.000
Human Social Support 1.189 2.724 1.009 4.287 3.517
@@
Table 9: F-value for self-report and human-perceived personality scores between high and low personality settings.
Note: NA refers to no within-group variance.
@@
Table 9 presents the F-values for human-perceived personality scores under high and low personality settings across five tasks. These values indicate the ratio of variances in personality perceptions between the two settings for each domain. The table shows that the variance between the high and low settings is more pronounced than within each group. Notably, this effect is most prominent in certain tasks and personality domains, such as Agreeableness and Conscientiousness.
Appendix H Statistical Significance and Correction
As shown in Table 10, the differences in chatbots’ self-reports between high and low conditions across tasks are both clear and statistically significant ($p<.001$). In contrast, while human-perceived scores also reflect differences between conditions, certain traits (e.g., Consciousness and Openness) show weaker or non-significant differences in specific contexts. This pattern suggests that changes in human perception are less pronounced and more sensitive to contextual factors than those captured by self-report measures.
Evaluation Domain Job Interview Public Service Social Support Travel Planning Guided Learning All Tasks
Human EXT .376.376 $001^{***}$ .287.287 $012^{*} .$404.404 $001^{***}$
AGR $^{***} .001.001^{**}$ $01^{***} .00$$2^{**} <.00$$1^{***} <.00$$01^{***}$
CON .428.428 $018^{*} .$942.942 $003^{**} .$$93^{*} .00$$04^{**}$
NEU $009^{**} .$$029^{*} .$256.256 $004^{**} .$$56^{*} <.00$$01^{***}$
OPE .539.539 $048^{*} .$$098^{*} .$$084^{*} .$737.737 $003^{**}$
BFI-2-XS EXT $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
AGR NA NA NA NA $^{***} <.001<.001^{***}$
CON NA NA $001^{***}$ NA NA $001^{***}$
NEU NA NA NA NA $^{***} <.001<.001^{***}$
OPE NA NA NA NA $^{***} <.001<.001^{***}$
BFI-2 EXT $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
AGR $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
CON $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
NEU $001^{***}$ NA $001^{***}$ NA $^{***} <.001<.001^{***}$
OPE $001^{***}$ NA $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
IPIP-NEO-120 EXT $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
AGR $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
CON $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
NEU $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
OPE $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$
Table 10: Comparing p-values of personality scores under high and low conditions.
Note: NA refers to no within-group variance. Comparisons within each domain are limited to chatbots explicitly prompted for that domain. A p-value <.05<.05 indicates statistical significance at the conventional threshold, while $p<.001$ denotes very strong evidence that the observed difference is unlikely to be due to chance.
Further evidence of this discrepancy is provided in Table 11, which presents results of paired t-tests comparing self-report and human-perceived personality scores within the same condition (either high or low). In the table, most of the p-values are highly significant in the three psychometric scales, further highlighting the differences between self-report and human-perceived personality scores. This mismatch supports the effectiveness of our prompt setting in distinguishing between static self-report metrics and real-world perceptions, and underscores the limitations of relying solely on self-report measures to evaluate the personality of personalized chatbots.
Evaluation Domain Job Interview Public Service Social Support Travel Planning Guided Learning All Tasks
High Low High Low High Low High Low High Low High Low
EXT $^{***} <.001<.001^{***}$ $05^{**} <.00$$1^{***} .00$$6^{**} <.00$$1^{***} .00$$3^{**} <.00$$1^{***} .00$$8^{**} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$
AGR $004^{**} .$182.182 $019^{*} .$$60^{*} .00$$9^{**} .00$$7^{**} .00$$02^{**}$ .101.101 $001^{**} .$$75^{*} <.00$$1^{***} <.00$$01^{***}$
BFI-2-XS CON $06^{**} <.00$$1^{***} .00$$9^{**} .00$$01^{**}$ $48^{*} <.00$$1^{***} .00$$5^{**} <.00$$01^{***}$ $24^{*} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$
NEU $01^{***} .00$$4^{**} <.00$$1^{***} <.00$$1^{***} <.00$$1^{***} .00$$8^{**} .00$$1^{**} .00$$9^{**} <.00$$1^{***} .00$$8^{**} <.00$$1^{***} <.00$$01^{***}$
OPE $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $14^{*} <.00$$01^{***}$ $22^{*} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$
EXT $^{***} <.001<.001^{***}$ .343.343 $01^{***} .00$$6^{**} <.00$$1^{***} .00$$4^{**} <.00$$01^{***}$ $41^{*} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$
AGR $039^{*} .$182.182 $019^{*} .$$063^{*} .$$010^{*} .$$^{*} .002.002^{**}$ .100.100 $014^{*} .$613.613 $^{***} <.001<.001^{***}$
BFI-2 CON .662.662 $03^{**} .00$$9^{**} .00$$01^{**}$ .409.409 $01^{***} .00$$5^{**} <.00$$01^{***}$ $36^{*} <.00$$1^{***} .00$$7^{**} <.00$$01^{***}$
NEU $001^{***}$ $12^{**} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$ .318.318 $01^{**} .00$$9^{**} <.00$$1^{***} .00$$9^{**} <.00$$1^{***} <.00$$01^{***}$
OPE $02^{**} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$ .116.116 $001^{***}$ $22^{*} <.00$$1^{***} .00$$7^{**} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$
EXT $^{***} <.001<.001^{***}$ $18^{*} <.00$$1^{***} .00$$4^{**} <.00$$1^{***} .00$$2^{**} <.00$$1^{***} .00$$5^{**} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$
AGR $06^{**} .00$$03^{**}$ $031^{*} .$910.910 $017^{*} .$$36^{*} .00$$02^{**}$ .115.115 $005^{**} .$$13^{*} <.00$$1^{***} .00$$02^{**}$
IPIP-NEO-120 CON $11^{*} <.00$$^{***} .011.0$$1^{*}$ $01^{**} .00$$9^{**} <.00$$1^{***} .00$$5^{**} <.00$$01^{***}$ $30^{*} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$
NEU $01^{***} .00$$6^{**} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$ $25^{*} .00$$02^{**}$ $36^{*} .00$$^{**} .011.0$$1^{*}$ $^{***} <.001<.001^{***}$
OPE $^{***} .010.010^{*}$ $01^{***} .00$$5^{**} <.00$$1^{***} .00$$06^{**}$ $15^{*} <.00$$1^{***} .00$$5^{**} .00$$3^{**} <.00$$1^{***} <.00$$01^{***}$
Table 11: Comparing p-values of self-report and human-perceived personality scores.
Note: Comparisons within each domain are limited to chatbots explicitly prompted for that domain. A p-value <.05<.05 indicates statistical significance at the conventional threshold, while $p<.001$ denotes very strong evidence that the observed difference is unlikely to be due to chance.
With the application of significance correction, our overall conclusions remain unaffected (Table 12). As we noted, most of our significant p-values in Table 11 are below .001.001. This threshold is more stringent than the adjusted alpha level that would result from correcting for 50 comparisons using either Bonferroni correction ($\alpha=.001$) or FDR procedures such as Benjamini-Hochberg. Thus, even under false discovery rate control (e.g., FDRFDR $q<0.05$), our main effects would still be retained as statistically significant. In this context, applying FWER/FDR correction would not alter the statistical interpretation of our key results, and would not undermine the broader theoretical conclusions drawn from them.
Evaluation Domain Job Interview Public Service Social Support Travel Planning Guided Learning
High Low High Low High Low High Low High Low
EXT $^{***} <.001<.001^{***}$ $08^{**} <.00$$1^{***} .00$$9^{**} <.00$$1^{***} .00$$5^{**} .00$$^{**} .011.0$$1^{*}$ $001^{***}$
AGR $006^{**} .$182.182 $022^{*} .$$064^{*} .$$011^{*} .$$10^{*} .00$$03^{**}$ .103.103 $002^{**} .$$078^{*}$
BFI-2-XS CON $09^{**} <.00$$^{***} .011.0$$1^{*}$ $002^{**} .$$52^{*} <.00$$1^{***} .00$$8^{**} <.00$$01^{***}$ $27^{*} <.00$$01^{***}$
NEU $01^{***} .00$$6^{**} <.00$$1^{***} <.00$$1^{***} <.00$$^{***} .011.0$$1^{*}$ $002^{**} .$$11^{*} <.00$$^{***} .011.0$$1^{*}$
OPE $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $^{***} <.001<.001^{***}$ $17^{*} <.00$$01^{***}$ $25^{*} <.00$$01^{***}$
EXT $02^{**} <.00$$01^{***}$ .363.363 $001^{***}$ $11^{*} <.00$$1^{***} .00$$7^{**} .00$$01^{**}$ $50^{*} <.00$$01^{***}$
AGR $049^{*} .$204.204 $027^{*} .$$076^{*} .$$015^{*} .$$27^{*} .00$$04^{**}$ .118.118 $020^{*} .$624.624
BFI-2 CON .662.662 $005^{**} .$$14^{*} .00$$02^{**}$ .424.424 $01^{***} .00$$9^{**} <.00$$01^{***}$ $46^{*} <.00$$01^{***}$
NEU $001^{***}$ $19^{*} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$ .343.343 $003^{**} .$$14^{*} .00$$01^{**}$ $014^{*}$
OPE $05^{**} <.00$$1^{***} <.00$$1^{***} <.00$$01^{***}$ .133.133 $001^{***}$ $28^{*} <.00$$^{***} .011.0$$1^{*}$ $001^{***}$
EXT $02^{**} <.00$$01^{***}$ $21^{*} <.00$$1^{***} .00$$7^{**} <.00$$1^{***} .00$$5^{**} .00$$1^{**} .00$$8^{**} <.00$$01^{***}$
AGR $09^{**} .00$$07^{**}$ $033^{*} .$910.910 $020^{*} .$$38^{*} .00$$05^{**}$ .118.118 $008^{**} .$$016^{*}$
IPIP-NEO-120 CON $14^{*} .00$$01^{**}$ $14^{*} .00$$04^{**}$ $12^{*} <.00$$1^{***} .00$$8^{**} <.00$$01^{***}$ $33^{*} <.00$$01^{***}$
NEU $01^{**} .00$$8^{**} <.00$$1^{***} .00$$1^{**} .00$$01^{**}$ $28^{*} .00$$04^{**}$ $38^{*} .00$$02^{**}$ $014^{*}$
OPE $001^{***}$ $14^{*} <.00$$1^{***} .00$$8^{**} <.00$$1^{***} .00$$08^{**}$ $19^{*} <.00$$1^{***} .00$$8^{**} .00$$06^{**}$
Table 12: FDR-adjusted p-values, referred to as q-values, calculated using the Benjamini-Hochberg procedure to compare self-report and human-perceived personality scores under high and low settings. Comparisons within each domain are limited to chatbots explicitly prompted for that domain.
Note: Comparisons within each domain are limited to chatbots explicitly prompted for that domain. A p-value <.05<.05 indicates statistical significance at the conventional threshold, while $p<.001$ denotes very strong evidence that the observed difference is unlikely to be due to chance.
To account for individual variability, we aggregate evaluation scores across chatbots that share the same personality configuration within each task setting (i.e., all high or all low conditions). Although these aggregated results do not account for finer distinctions between specific tasks, they do consider the effects of high- or low-condition personality settings. As shown in Table 13, human perception scores are generally lower than self-report scores under high personality settings and higher under low settings, with confidence intervals that do not include zero. This suggests that personality traits configured in the chatbot are perceived by users to a lesser extent than intended in most cases. Moreover, the differences between self-report and human-perceived scores are statistically significant across all five personality domains, underscoring a notable discrepancy between chatbot self-reports and human perceptions at two personality settings.
Evaluation Domain High Low
Mean CI p-value Mean CI p-value
EXT -0.90 (-1.10, -0.70) $001^{***}$ 2.27 (2.02, 2.53) $001^{***}$
AGR -0.89 (-1.14, -0.65) $001^{***}$ 0.67 (0.37, 0.97) $001^{***}$
BFI-2-XS CON -0.81 (-1.07, -0.56) $001^{***}$ 2.54 (2.21, 2.87) $001^{***}$
NEU -2.23 (-2.55, -1.92) $001^{***}$ 0.79 (0.60, 0.98) $001^{***}$
OPE -1.20 (-1.43, -0.97) $001^{***}$ 2.27 (2.05, 2.50) $001^{***}$
EXT -0.80 (-1.05, -0.55) $001^{***}$ 2.19 (1.93, 2.45) $001^{***}$
AGR -0.59 (-0.87, -0.30) $001^{***}$ 0.59 (0.28, 0.90) $001^{***}$
BFI-2 CON -0.39 (-0.67, -0.11) $007^{**}$ 2.31 (1.98, 2.65) $001^{***}$
NEU -2.12 (-2.44, -1.79) $001^{***}$ 0.69 (0.47, 0.91) $001^{***}$
OPE -1.00 (-1.26, -0.74) $001^{***}$ 2.23 (2.01, 2.46) $001^{***}$
EXT -0.89 (-1.09, -0.68) $001^{***}$ 1.71 (1.47, 1.95) $001^{***}$
AGR -0.80 (-1.04, -0.56) $001^{***} -$0.68 (-1.11, -0.25) .002.002
IPIP-NEO-120 CON -0.82 (-1.06, -0.57) $001^{***}$ 2.34 (2.01, 2.66) $001^{***}$
NEU -1.97 (-2.28, -1.66) $001^{***}$ 0.69 (0.50, 0.89) $001^{***}$
OPE -1.18 (-1.40, -0.96) $001^{***}$ 1.33 (1.04, 1.62) $001^{***}$
Table 13: Mean differences, 95% confidence intervals (CIs), and p-values between human-perceived and self-report personality scores under high and low personality settings (computed as human evaluation minus self-report).
Note: A p-value <.05<.05 indicates statistical significance at the conventional threshold, while $p<.001$ denotes very strong evidence that the observed difference is unlikely to be due to chance.
Appendix I Correlation Analysis
Table 14 presents the correlation analysis between self-report and human-perceived personality scores using the BFI-2-XS, BFI-2, and IPIP-NEO-120 questionnaires. While Agreeableness consistently demonstrates moderate positive correlations, Conscientiousness shows weak or negative correlations in several cases, such as -0.01 in the Social Support task with BFI-2-XS. Extraversion also shows near-zero correlations in the social support task,, with values of 0.08 (BFI-2-XS), 0.02 (BFI-2), and -0.03 (IPIP-NEO-120). These results show a weak correlation between human-perceived personality and those evaluated via standard tests.
Evaluation Task EXT AGR CON NEU OPE
@@ Appendix I Correlation Analysis
Public Service 0.42 0.58 0.46 0.26 0.26
BFI-2-XS Social Support 0.08 0.63 -0.01 0.09 0.19
@@
Public Service 0.40 0.59 0.50 0.32 0.25
BFI-2 Social Support 0.02 0.64 -0.01 0.11 0.14
@@
Guided Learning 0.05 0.54 -0.03 0.30 0.08
@@
Public Service 0.37 0.57 0.45 0.33 0.25
IPIP-NEO-120 Social Support -0.03 0.48 -0.07 0.18 0.12
@@
Table 14: Correlation analysis between self-report and human-perceived personality scores.
Appendix J Fine-tuning Details
For fine-tuning, we split all the scripts of collected human-chatbot conversations and their corresponding human evaluations of chatbot personality into training and testing sets using an 80:20 ratio. Human-perceived chatbot personality scores were collected using the BFI-XS scale, which balances efficiency with robust predictive validity. We then fine-tuned GPT-4o (OpenAI, 2024) on the training set with instructions on how to analyze human-chatbot conversational transcripts and rate BFI-XS statements based on chatbot responses. Table 15 details the specific instructions used, where transcript refers to the human-chatbot conversational scripts. The evaluation was conducted on the testing set to assess the alignment between its output and actual human perceptions.
As shown in Table 16, the correlation between machine-inferred and human-perceived scores on the test set is higher than that of self-report scores across all scales. This finding suggests that personality evaluations based on multi-turn interactions can reduce their gap with human perceptions.
You are an AI assistant specializing in text analysis. Your task is to evaluate the personality of a chatbot based on its responses in a multi-turn conversation, focusing on the Big Five Personality traits. You will receive a transcript of the chatbot-user conversation and personality statements about the chatbot.
Analyzing the conversation transcript and considering the chatbot’s responses, indicate the extent to which you agree or disagree with the statement on a scale from 1 to 5 (where 1 = “disagree strongly”, 2 = “disagree a little”, 3 = “neither agree nor disagree”, 4 = ‘”agree a little”, and 5 = “agree strongly”).
Conversation Transcript: {transcript}
Table 15: Instructions for fine-tuning the model on human-chatbot conversation scripts.
Evaluation EXT AGR CON NEU OPE
BFI-2-XS 0.23 0.59 0.24 0.29 0.24
BFI-2 0.20 0.59 0.25 0.33 0.21
IPIP-NEO-120 0.17 0.56 0.25 0.36 0.24
Machine-inferred 0.28 0.77 0.52 0.40 0.32
Table 16: Correlation of self-report and machine-inferred personality scores with human-perceived scores.
Appendix K Future Direction
This paper primarily examines, at the trait level, the relationship between the chatbot’s self-report scores and human-perceived scores, including correlations between trait scores, analyses of variance, and correlations between traits and external variables. In future research, it would be valuable to further explore differences in response patterns between chatbot self-report scores and human-perceived scores (e.g., Sun et al., 2019); or to investigate how the two differ when responding to forced-choice scales (e.g., Zhang et al., 2024).
Also, the key point mentioned in this paper (i.e., measuring effectiveness in application scenarios) is not limited to text-based LLMs. They can be extended to a broader range of models, such as MLLMs or VLMs, which operate across multiple input modalities and are increasingly applied to complex real-world tasks. It is also necessary to conduct more measurements from the human perspective for these models (e.g., Li et al., 2024).
Overall, the evaluation of chatbots or LLMs can draw extensively on well-established psychometric theories and frameworks from the field of psychology (e.g., Jiang et al., 2025), such as classical test theory (CTT) and item response theory (IRT), which can provide the theoretical foundation for developing more systematic and precise evaluation methods.
Appendix L Limitations
Several limitations exist in the current study. First, there may be bias in the choice of psychometric tests. Although we sought to minimize this by using personality assessments of different lengths for self-report data, the human-perceived personality was measured with only one questionnaire due to time constraints. Future research could address this by incorporating a broader range of personality tests and developing assessments specifically tailored for LLMs to achieve more accurate measurements.
Second, our findings are based on a single chatbot personality setting method (i.e., prompt-based control) and may not generalize well to others, highlighting the need for further investigation with different approaches. Additionally, the chatbot was designed with a focus on task completion, mirroring real-world applications to ensure ecological validity. As a result, the task settings may have constrained the variability in perceived personality, especially if the personality traits conflict with the intended function or design goals of the chatbot. For instance, human-perceived Neuroticism scores remained relatively low across both high and low conditions, suggesting that LLMs may implicitly suppress negative or anxious traits during actual task performance, even when explicitly prompted to adopt a neurotic persona. That indicates a potential limitation of prompt-based personality design. Moreover, the evaluation was conducted on five common chatbot tasks, which may not capture the full spectrum of user interactions. Expanding the task set in future studies could provide a more comprehensive view of task-based chatbot personality assessment.
Third, while our study only focuses on GPT-4o (OpenAI, 2024), we expect the same experimental setup could be extended to other LLMs. The extent to which our findings generalize to a broader range of models remains an open question for future research. Furthermore, while LLMs perform well on benchmarks across multiple languages, this study evaluated the model using English psychometric tests and English-speaking participants. Future research could explore personality dimensions within culture-specific contexts to provide more diverse insights on chatbot design.